# Roadmaps with Gaps over Controllers: Achieving Efficiency in Planning under Dynamics

### v3

@@ Abstract
This paper aims to improve the computational efficiency of motion planning for mobile robots with non-trivial dynamics through the use of learned controllers. It adopts a decoupled strategy, where a system-specific controller is first trained offline in an empty environment to deal with the robot’s dynamics. For a target environment, the proposed approach constructs offline a data structure, a “Roadmap with Gaps,” to approximately learn how to solve planning queries in this environment using the learned controller. The nodes of the roadmap correspond to local regions. Edges correspond to applications of the learned control policy that approximately connect these regions. Gaps arise because the controller does not perfectly connect pairs of individual states along edges. Online, given a query, a tree sampling-based motion planner uses the roadmap so that the tree’s expansion is informed towards the goal region. The tree expansion selects local subgoals given a wavefront on the roadmap that guides towards the goal. When the controller cannot reach a subgoal region, the planner resorts to random exploration to maintain probabilistic completeness and asymptotic optimality. The accompanying experimental evaluation shows that the approach significantly improves the computational efficiency of motion planning on various benchmarks, including physics-based vehicular models on uneven and varying friction terrains as well as a quadrotor under air pressure effects. Website: https://prx-kinodynamic.github.io/projects/rogue
@@ I Introduction
Kinodynamic motion planning is useful for mobile robots with non-trivial dynamics that must negotiate environments with obstacles, such as a warehouse, or physical features like uneven terrain and ice. The problem is challenging when there is no access to a local steering function that can connect two robot states. Tree sampling-based kinodynamic planners [1] do not need a steering function as they only forward propagate the system’s dynamics. Recent variants of these planners provide asymptotic optimality (AO) via the propagation of randomly sampled controls at every iteration [2, 3, 4, 5]. The use of random controls, however, while important for theoretical properties, often results in slow convergence to high-quality solutions in practice.
This work aims to improve the efficiency of such planners using machine learning. Deep Reinforcement Learning (DRL) is a promising direction to improve the practical convergence of motion planners [6]. DRL can learn a local controller that approximately steers the robot between two states without collision. It can then be integrated inside a planner like a Probabilistic Roadmap (PRM) [7] or a Rapidly-exploring Random Tree (RRT) [8]. DRL is applicable in kinodynamic motion planning unlike some supervised learning alternatives [9, 10, 11, 12], which require a steering function. DRL suffers, however, from large training data requirements and requires proper tuning of reward parameters [13].
Most current DRL solutions also lack long-horizon planning capabilities, meaning that a planner is still needed for the global search of the state space. Several recent works have leveraged DRL to build an abstract representation of the planning problem and search for an optimal path between the start and the goal on this representation. Search on the Replay Buffer (SoRB) [14] and Sparse Graphical Memory (SGM) [15] build a graph where the nodes correspond to states visited by the DRL agent during training. Similar to SoRB and SGM, Deep Skill Graphs [16, 17, 18] builds a graph representation, where nodes correspond to subgoals, and edges correspond to control policies between them. A common feature of these methods is that the control policy is trained in the planning environment of interest. This requires the learning algorithm to reason jointly about the system’s dynamics and the obstacles present in the environment, which is challenging for second-order systems [19].
@@
Figure 1: Solution trajectories (thick lines) and planning trees (thin lines) for a MuSHR vehicle over an uneven terrain simulated in MuJoCo. The cost function is trajectory duration, which is impacted by the uneven terrain. The proposed expansion method for kinodynamic planning (left) leverages a roadmap with gaps to avoid difficult terrain resulting in shorter duration solutions. The alternative expansion (right), which samples random local goals, navigates the rough hills less effectively. It results in a shorter distance path but results in a much slower trajectory. Planning time was 60s for both methods for the computed trajectories.
@@
Prior work integrates DRL in kinodynamic motion planning via a decoupled strategy [20, 21], where a controller is first trained offline in an empty environment and the learning process only deals with the system’s dynamics. The absence of obstacles allows learning the controller with significantly fewer data and only a sparse reward. During online planning, these prior efforts engineer local goals that bias expansion toward the global planning goal via manual engineering. The local goals are passed to the controller, which outputs a control towards them. Designing, however, an informed local goal procedure requires manual effort, which is undesirable and may not apply equally well across environments. Furthermore, the local goals these procedures generate are unaware of the trained controller’s reachability properties.
@@
This work proposes a data structure, a Roadmap with Gaps, that learns the approximate reachability of local regions in a given environment given a controller. The approach constructs a directed graph in the free configuration space of the robot. An edge exists from a source node/region to a target when the source node’s nominal state can reach the target node’s termination set without collisions by application of the controller. Gaps arise since the learned controller does not connect any two exact states on this roadmap perfectly. Given a new planning problem in this environment, a wavefront expressing the cost-to-goal on this roadmap provides every node with the direction of motion towards the goal. This roadmap guidance is integrated with an AO tree sampling-based planner, where the accessible neighboring roadmap node with the lowest cost towards the global goal is selected as the expansion target for the tree planner.
@@
The proposed approach provides a tradeoff. It incurs an offline computational cost for pre-processing a known environment and the robot’s dynamics to provide online computational benefits for multiple planning queries in the same workspace. At the same time it retains AO properties and is an informed, efficient solution by understanding the reachability properties of the control method used for the target system and the environment’s topology. Prior approaches that guide tree sampling-based planners using graphical abstractions [22, 23] do not provide desirable properties (e.g., AO) or do not reason about the system’s dynamics. A related approach proposed Bundle of Edges (BoE) for kinodynamic planning [24] but relies on random control propagation, which reduces efficiency.
In summary, the paper’s contributions are the following:
@@
  • It introduces the “Roadmap with Gaps,” a data structure that provides high-level guidance how a robot can navigate a target environment given a (potentially learned) controller for the target robotic system.
@@
  • It adapts an existing framework for Asymptotically Optimal (AO) kinodynamic motion planning to utilize the “Roadmap with Gaps” and achieve improved efficiency. This results in the proposed Roadmap-Guided Expansion (RoGuE)-Tree method.
@@
The accompanying evaluation demonstrates the effectiveness of RoGuE both on analytically and physically simulated mobile robots in various environments, including physics-based vehicular models on uneven and varying friction terrains as well as a quadrotor under air pressure effects. RoGuE results in lower-cost solutions faster than random propagation given an AO sampling-based kinodynamic planner.
@@ III Proposed Method
Overall, the proposed method has 2 offline stages: (i) training a controller $\pi(x,q)$ in an obstacle-free environment, and (ii) building a “Roadmap with Gaps” $(\mathcal{V,E})$ in the target environment given the available controller. Finally, it has an online phase given a motion planning query $(x_{0},x_{G})$ in the target environment. Online, a sampling-based kinodynamic planner expands a tree of feasible trajectories guided by a wavefront function computed over the “Roadmap with Gaps”.
@@ III-B Building a Roadmap with Gaps offline
The roadmap construction procedure (Fig 2(a)) samples first a set of configuration milestones $\{q_{i}\}_{i=1}^{|N|}$, which form the vertices $\mathcal{V}$ in the roadmap. In this work, the milestones are chosen over a grid in $\mathbb{Q}$. They are collision-checked and verified to be in $\mathbb{Q}_{\mathrm{f}}$. Then, for the generation of edges $\mathcal{E}$, the procedure selects a random configuration $q_{i}\in\mathcal{V}$ and obtains two sets: $A$ and $D$. $A$ is the set of all vertices $q\in\mathcal{V},$ whose hyperballs $\mathcal{B}(q,\epsilon)$ can be reached from $x_{i}=\mathbb{M}^{-1}(q_{i})$ given the controller $\pi$. Similarly, $D$ is the set of vertices $q\in\mathcal{V}$, where the state $x=\mathbb{M}^{-1}(q)$ can reach the hyperball $\mathcal{B}(q_{i},\epsilon)$ given the controller $\pi$. Then, edges from $q_{i}$ to vertices in $A$ and edges from vertices in $D$ to $q_{i}$ are added to the set $\mathcal{E}$.
The application of the controller for the edge construction, especially a learned one for a non-linear dynamical system, implies that for all edges $(q_{i},q_{j})$ in the roadmap, there is no guarantee that the resulting configuration $q_{j}$ can be achieved exactly from $q_{i}$. Instead, the corresponding edge only guarantees that the system will be within a hyperball $\mathcal{B}(q_{G},\epsilon)$ if it is initialized at $\mathbb{M}^{-1}(q_{i})$. Consequently, the edges of the roadmap introduce “gaps” as highlighted by the figure. This issue is exacerbated if the controller is used to follow a sequence of edges on the roadmap, e.g., $(q_{i},q_{j})$ and $(q_{j},q_{k})$. Since the controller’s application over the first edge can only bring the system in the vicinity of $q_{j}$, the application of the controller corresponding to the second edge at the resulting configuration may not even bring the system within an $\epsilon$-hyperball of $q_{k}$. Consequently, paths on the “Roadmap with Gaps” do not respect the dynamic constraints of the vehicle and cannot be used to directly solve kinodynamic planning problems. They can still provide, however, useful guidance for a kinodynamic planner.
Figure 2: (a) A Roadmap with Gaps consists of vertices (configurations, represented as dotted boxes). The roadmap’s directional edges (dotted lines) correspond to where the controller was successfully executed from the source to the target configuration, given some tolerance. (b) For a new planning query $(x_{0},x_{G})$, the start $q_{0}$ and goal $q_{G}$ are added to the roadmap (transparent). Given the current planning tree (opaque), Roadmap-Guided Expansion (RoGuE) selects informed local goals for the controller $\pi$. The closest roadmap node $q_{\text{near}}$ to the selected tree node $x_{\text{sel}}$ is identified. Its successor is passed as the local goal $q_{\text{lg}}$ to the controller. This expansion adds a new tree node $x_{\text{new}}$.
@@ 9             EXTEND-TREE($\mathrm{T},x_{\text{sel}}\rightarrow x_{\text{new}}$);
RoGuE (Algo. 3, Fig 2(b)) is an expansion procedure that uses the roadmap wavefront information $\mathcal{W}$ to provide an informed local goal to $\pi$. The first time a tree node $x_{\text{sel}}$ is selected for expansion, RoGuE identifies the closest roadmap node $q_{\text{near}}\in\mathcal{V}$ according to the distance function $d(\cdot,\cdot$) (Line 2). It then queries the Successor($q_{\text{near}}$) given the wavefront $\mathcal{W}$. If ($q_{\text{near}}$) is defined, it is provided as the local goal $q_{\text{lg}}$ to the controller $\pi$ (Lines 3-5). If the successor is not defined, a random local goal in $\mathbb{Q}$ is provided (Line 6-7).
@@ IV Experimental Evaluation
Figure 3: Physically simulated benchmarks using MuJoCo. (L-R) Maze, Friction, Quadrotor.
@@
Figure 4: Evaluation on the analytically simulated benchmarks. Left and middle-left: higher success rate earlier is better. Right and middle-right: lower path cost is better. As more solutions are discovered, additional solutions for harder problems are discovered, so in some cases, the average cost may increase over time. Each problem instance is run 10 times to account for different random seeds.
@@ IV-A Results on analytically simulated benchmarks
Across all benchmarks, for the differential drive and car-like dynamics, DIRT-RoGuE finds the lowest cost solutions overall. It also is the only expansion strategy that returns solutions in all trials. Among the RRT expansion strategies, RRT-RoGuE achieves the highest success rate across all benchmarks and the lowest cost solutions overall. RRT-RoGuE also achieves comparable performance (both in terms of success rate and solution quality) to the informed DIRT-Random an DIRT-RLG planners.
@@ IV-B Results on physically simulated benchmarks
Figs. 1 and  3 visualize the different environments in MuJoCo, and Fig. 5 provides the experimental results. Only the DIRT motion planner is reported in these experiments as the RRT-based solutions cannot find a solution within the allotted times. Since each call to the MuJoCo engine is expensive, all expansion strategies use a blossom $k=1$.

### v5

@@ Abstract
This paper aims to improve the computational efficiency of motion planning for mobile robots with non-trivial dynamics through the use of learned controllers. Offline, a system-specific controller is first trained in an empty environment. Then, for the target environment, the approach constructs a data structure, a “Roadmap with Gaps,” to approximately learn how to solve planning queries using the learned controller. The roadmap nodes correspond to local regions. Edges correspond to applications of the learned controller that approximately connect these regions. Gaps arise as the controller does not perfectly connect pairs of individual states along edges. Online, given a query, a tree sampling-based motion planner uses the roadmap so that the tree’s expansion is informed towards the goal region. The tree expansion selects local subgoals given a wavefront on the roadmap that guides towards the goal. When the controller cannot reach a subgoal region, the planner resorts to random exploration to maintain probabilistic completeness and asymptotic optimality. The accompanying experimental evaluation shows that the approach significantly improves the computational efficiency of motion planning on various benchmarks, including physics-based vehicular models on uneven and varying friction terrains as well as a quadrotor under air pressure effects. Website: https://prx-kinodynamic.github.io/projects/rogue
@@ I Introduction
Kinodynamic motion planning allows mobile robots with non-trivial dynamics to negotiate environments with obstacles, such as a warehouse, or physical features, such as uneven terrain and ice. The problem is challenging when there is no local steering function available that connects two states. Tree sampling-based planners [1] do not need a steering function as they only forward propagate the system’s dynamics. Some variants provide Asymptotic Optimality (AO) by propagating randomly sampled controls [2, 3, 4, 5]. While random controls provide theoretical properties, they result in slow convergence to high-quality solutions in practice.
This work aims to improve the efficiency of such AI kinodynamic planners [6, 7] by integrating controllers trained via Reinforcement Learning (RL) [8]. Supervision has also be used to train controllers that are then integrated with planners [9, 10, 11, 12]. This is often done, however, under the assumption that a steering function is available for the robot, which is not assumed here. While one may attempt to solve planning problems directly with RL, this often suffers from large training data requirements and requires careful tuning of rewards [13]. Most RL solutions lack long-horizon reasoning, which a planner provides.
Several recent works have leveraged RL to build an abstract representation of the planning problem and search for an optimal path between the start and the goal. Search on the Replay Buffer (SoRB) [14] and Sparse Graphical Memory (SGM) [15] build a graph where the nodes correspond to states visited by the RL agent during training. Similar to SoRB and SGM, Deep Skill Graphs [16, 17, 18] builds a graph representation, where nodes correspond to subgoals, and edges correspond to control policies between them. A common feature of these methods is that the control policy is trained in the planning environment of interest. This requires the learning algorithm to reason jointly about the system’s dynamics and the obstacles present in the environment, which is challenging for second-order systems [19].
@@
Figure 1: Solution trajectories (thick lines) and planning trees (thin lines) for a MuSHR vehicle over an uneven terrain in MuJoCo. The cost function is trajectory duration, which is impacted by the uneven terrain. The proposed expansion method for kinodynamic planning (left) leverages a “Roadmap with Gaps” to avoid difficult terrain resulting in shorter duration solutions. The alternative expansion (right), which samples random local goals, navigates the rough hills less effectively. It results in a shorter distance path but a much slower trajectory. Planning time was 60s for both methods.
@@
Prior work by the authors integrates RL in kinodynamic motion planning via a decoupled strategy [20, 21], where a controller is first trained offline in an empty environment and the learning process only deals with the system’s dynamics. The absence of obstacles allows learning the controller with significantly fewer data and only a sparse reward. During online planning, these prior efforts engineer local goals that bias expansion toward the global goal via manual engineering. Designing, however, an informed local goal procedure requires manual effort, which is undesirable and may not work well across environments. Furthermore, the local goals these procedures generate are unaware of the trained controller’s reachability properties.
@@
This work proposes a data structure, a Roadmap with Gaps, that learns the approximate reachability of local regions in a given environment given a controller. The approach constructs a directed graph in the free configuration space of the robot. An edge exists from a source node/region to a target when the source node’s nominal state can reach the target node’s termination set without collisions by applying the controller. Gaps arise since the learned controller does not connect any two exact states exactly. Given a new planning query, a wavefront expressing the cost-to-goal on this roadmap provides every node with the direction of motion towards the goal. This roadmap guidance is integrated with an AO tree sampling-based planner. During the tree planner’s expansion, nodes that are closer to the global goal given the roadmap guidance are prioritized.
@@
The proposed approach provides a tradeoff. It incurs an offline computational cost for pre-processing a known environment and the robot’s dynamics to provide online computational benefits for multiple planning queries in the same workspace. At the same time it retains AO properties and is an informed, efficient solution by understanding the reachability properties of the learned controller and the environment’s connectivity. Prior approaches that guide tree sampling-based planners using graphical abstractions [22, 23] do not provide desirable properties (e.g., AO) or do not reason about the system’s dynamics. A related kinodynamic planning approach introduced the Bundle of Edges (BoE) notion [24] but relies on random control propagation, which reduces efficiency.
In summary, this paper proposes the Roadmap-Guided Expansion (RoGuE)-Tree method, which:
  • $\circ$
@@
    Introduces the “Roadmap with Gaps,” a data structure that approximately guides how a robot can navigate a target environment given a controller, and
  • $\circ$
@@
    Adapts an Asymptotically Optimal (AO) framework for kinodynamic planning to utilize the “Roadmap with Gaps”.
@@
The accompanying evaluation shows that RoGuE results in lower-cost solutions faster than random propagation given an AO sampling-based kinodynamic planner both on analytically and physically simulated models in various environments. This includes physics-based vehicular models on uneven and varying friction terrains as well as a quadrotor under air pressure effects.
@@ III Proposed Method
Figure 2: Stages of the proposed pipeline. Offline: A controller is trained in an empty environment. The roadmap with gaps is built over the robot’s free C-space in a target environment. Online: Given a new query $(x_{0},x_{G})$, a wavefront is computed over the roadmap. At every iteration of the tree sampling-based planner, a local goal is computed given the wavefront information, and a candidate control is propagated towards the local goal via the trained controller.
@@
Overall, the proposed method (Fig 2) has 2 offline stages: (i) training a controller $\pi(x,q)$ in an obstacle-free environment, and (ii) building a “Roadmap with Gaps” $(\mathcal{V,E})$ in the target environment given the available controller. Finally, it has an online phase given a motion planning query $(x_{0},x_{G})$ in the target environment. Online, a sampling-based kinodynamic planner expands a tree of feasible trajectories guided by a wavefront function computed over the “Roadmap with Gaps”.
@@ III-B Building a Roadmap with Gaps offline
The roadmap construction procedure (Fig 3) samples first a set of configuration milestones $\{q_{i}\}_{i=1}^{|N|}$, which form the vertices $\mathcal{V}$ in the roadmap. In this work, the milestones are chosen over a grid in $\mathbb{Q}$. They are collision-checked and verified to be in $\mathbb{Q}_{\mathrm{f}}$. Then, for the generation of edges $\mathcal{E}$, the procedure selects a random configuration $q_{i}\in\mathcal{V}$ and obtains two sets: $A$ and $D$. $A$ is the set of all vertices $q\in\mathcal{V},$ whose hyperballs $\mathcal{B}(q,\epsilon)$ can be reached from $x_{i}=\mathbb{M}^{-1}(q_{i})$ given the controller $\pi$. Similarly, $D$ is the set of vertices $q\in\mathcal{V}$, where the state $x=\mathbb{M}^{-1}(q)$ can reach the hyperball $\mathcal{B}(q_{i},\epsilon)$ given the controller $\pi$. Then, edges from $q_{i}$ to vertices in $A$ and edges from vertices in $D$ to $q_{i}$ are added to the set $\mathcal{E}$.
The application of the controller for the edge construction, especially a learned one for a non-linear dynamical system, implies that for all edges $(q_{i},q_{j})$ in the roadmap, there is no guarantee that the resulting configuration $q_{j}$ can be achieved exactly from $q_{i}$. Instead, the corresponding edge only guarantees that the system will be within a hyperball $\mathcal{B}(q_{G},\epsilon)$ if it is initialized at $\mathbb{M}^{-1}(q_{i})$.
Consequently, the edges of the roadmap introduce “gaps” as highlighted by the figure. This issue is exacerbated if the controller is used to follow a sequence of edges on the roadmap, e.g., $(q_{i},q_{j})$ and $(q_{j},q_{k})$. Since the controller’s application over the first edge can only bring the system in the vicinity of $q_{j}$, the application of the controller corresponding to the second edge at the resulting configuration may not even bring the system within an $\epsilon$-hyperball of $q_{k}$. Consequently, paths on the “Roadmap with Gaps” do not respect the dynamic constraints of the vehicle and cannot be used to directly solve kinodynamic planning problems. They can still provide, however, useful guidance for a kinodynamic planner.
Figure 3: (a) A Roadmap with Gaps consists of vertices (configurations, represented as dotted boxes). The roadmap’s directional edges (dotted lines) correspond to where the controller was successfully executed from the source to the target configuration, given some tolerance. (b) For a new planning query $(x_{0},x_{G})$, the start $q_{0}$ and goal $q_{G}$ are added to the roadmap. (c) Due to gaps, naively following the roadmap’s shortest path to the goal may not be feasible.
@@ 9             EXTEND-TREE($\mathrm{T},x_{\text{sel}}\rightarrow x_{\text{new}}$);
RoGuE (Algo. 3, Fig 4) is an expansion procedure that uses the roadmap wavefront information $\mathcal{W}$ to provide an informed local goal to $\pi$. The first time a tree node $x_{\text{sel}}$ is selected for expansion, RoGuE identifies the closest roadmap node $q_{\text{near}}\in\mathcal{V}$ according to the distance function $d(\cdot,\cdot$) (Line 2). It then queries the Successor($q_{\text{near}}$) given the wavefront $\mathcal{W}$. If ($q_{\text{near}}$) is defined, it is provided as the local goal $q_{\text{lg}}$ to the controller $\pi$ (Lines 3-5). If the successor is not defined, a random local goal in $\mathbb{Q}$ is provided (Line 6-7).
@@ 5             $u\leftarrow\pi(x_{\text{sel}},q_{\text{lg}});$
Figure 4: Roadmap-Guided Expansion (RoGuE): Given the roadmap with gaps in an environment (transparent) and the current planning tree (opaque), RoGuE selects informed local goals for the controller $\pi$. The closest roadmap node $q_{\text{near}}$ to the selected tree node $x_{\text{sel}}$ is identified. Its successor is passed as the local goal $q_{\text{lg}}$ to the controller. This expansion adds a new tree node $x_{\text{new}}$.
@@ IV Experimental Evaluation
Figure 5: Physically simulated benchmarks using MuJoCo. Top: (L-R) Maze, Terrain. Bottom: (L-R) Friction, Quadrotor.
@@
(a) Planning results on the Narrow benchmark.
  Refer to caption Refer to caption
(b) Planning results on the Indoors benchmark.
  Refer to caption Refer to caption
(c) Planning results on the Warehouse benchmark.
Figure 6: Left and middle-left: higher success rate earlier is better. Right and middle-right: lower path cost is better. As more solutions are discovered, additional solutions for harder problems are discovered, so in some cases, the average cost may increase over time. Each problem instance is run 10 times to account for different random seeds.
@@ IV-A Results on analytically simulated benchmarks
Across all benchmarks, for the differential drive and car-like dynamics, DIRT-RoGuE finds the lowest cost solutions overall. It also is the only expansion strategy that returns solutions in all trials. Among the RRT expansion strategies, RRT-RoGuE achieves the highest success rate across all benchmarks and the lowest cost solutions overall. RRT-RoGuE also achieves comparable performance (both in terms of success rate and solution quality) to the informed DIRT-Random an DIRT-RLG planners. Both DIRT-RLG and DIRT-Random fail to find solutions in all trials on the Indoor and Warehouse benchmarks and are slow to return solutions relative to DIRT-RoGuE.
In the Narrow benchmark, the shortest duration trajectories must traverse all the narrow passages. So although DIRT-Random and DIRT-RLG find solutions in all trials for the second-order differential drive, DIRT-RoGuE returns significantly lower-cost solutions much earlier.
@@ IV-B Results on physically simulated benchmarks
Fig. 5 visualize the different environments in MuJoCo, and Fig. 7 provides the experimental results. Only the DIRT motion planner is reported in these experiments as the RRT-based solutions cannot find a solution within the allotted times. Since each call to the MuJoCo engine is expensive, all expansion strategies use a blossom $k=1$.
@@
IV-C Ablation studies
Figure 8: Ablation experiments for various roadmap sizes.
Fig. 8 considers 3 roadmaps of different sizes by varying the number of configurations $N$ during roadmap construction: Baseline ($|\mathcal{V}|=1088,|\mathcal{E}|=10911$), Denser ($|\mathcal{V}|=1360,|\mathcal{E}|=156936$), and Sparser ($|\mathcal{V}|=688,|\mathcal{E}|=4718$). Planners using the Baseline and Denser roadmaps find competitive solutions quickly, while the Sparser roadmap cannot do so, motivating the use of large roadmaps for RoGuE.
Figure 9: Ablation experiments varying the max. edge duration.
Fig. 9 studies the effect of the maximum allowed edge cost $T_{\text{max}}$ on the online planner. Four different values of $T_{\text{max}}$ are considered: Baseline (10s), Shorter (5s), Longer (20s), and Longest (30s). The planner using the Shorter roadmap underperforms, while the planners using the other values find solutions quickly. As $T_{\text{max}}$ increases, the time taken by the planner to find solutions to all problems increases slightly without significantly affecting the cost of the returned solution. This suggests that while longer edges may help find high-quality solutions, they may also suffer due to the gaps in the roadmap.